OpenStudy (anonymous):

Please help me solve A developer wants to enclose a rectangular grassy lot that borders a city street for parking. If the developer has 272ft of fencing and does not fence the side along the street, what is the largest area that can be enclosed?

OpenStudy (anonymous):

2w + l = 272 ft ----> l = -2w + 272 => A = w ( -2w + 272) = - 2w² + 272w The largest area happens when A' = 0 A' = - 4w + 272 = 0 ---> w = 272/4 = 68 ft ==> l = -2* 68 + 272 = 136 ft Thus largest area can be enclosed: A = 68 * 136 = 9248 ft²

OpenStudy (anonymous):

THANKS